# Understanding Bias Reinforcement in LLM Agents Debate

### v4

@@ Abstract
Large Language Models (LLMs) solve complex problems using training-free methods like prompt engineering and in-context learning, yet ensuring reasoning correctness remains challenging. While self-correction methods such as self-consistency and self-refinement aim to improve reliability, they often reinforce biases due to the lack of effective feedback mechanisms. Multi-Agent Debate (MAD) has emerged as an alternative, but we identify two key limitations: bias reinforcement, where debate amplifies model biases instead of correcting them, and lack of perspective diversity, as all agents share the same model and reasoning patterns, limiting true debate effectiveness. To systematically evaluate these issues, we introduce MetaNIM Arena, a benchmark designed to assess LLMs in adversarial strategic decision-making, where dynamic interactions influence optimal decisions. To overcome MAD’s limitations, we propose DReaMAD (Diverse Reasoning via Multi-Agent Debate with Refined Prompt), a novel framework that (1) refines LLMs’ strategic prior knowledge to improve reasoning quality and (2) promotes diverse viewpoints within a single model by systematically modifying prompts, reducing bias. Empirical results show that DReaMAD significantly improves decision accuracy, reasoning diversity, and bias mitigation across multiple strategic tasks, establishing it as a more effective approach for LLM-based decision-making.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable problem-solving capabilities across a wide range of tasks by leveraging knowledge acquired from vast datasets (Achiam et al., 2023; The, ; Team et al., 2023; Dubey et al., 2024). These models can address complex decision-making problems using training-free methods such as prompt engineering (White et al., 2023; Chen et al., 2023a; Schulhoff et al., 2024b) and in-context learning (Brown et al., 2020; Dong et al., 2022; Wei et al., 2022; Mavromatis et al., 2023; Pan, 2023), which provide guidance for effective reasoning. However, these approaches do not explicitly guarantee the correctness of the generated responses.
To address this, recent research has explored self-correction mechanisms that allow LLMs to refine their own outputs without external feedback. Self-consistency (Wang et al., 2022; Chen et al., 2023c) enhances reliability by ensembling multiple responses. Self-refinement (Wan et al., 2023; Shinn et al., 2023; Madaan et al., 2024) enables LLMs to iteratively critique and revise their outputs. However, self-consistency lacks a critical feedback mechanism, meaning it does not iteratively refine responses but merely reduces the model’s inherent randomness by converging on the most frequently generated answer. Further recent studies (Huang et al., 2024) suggest that self-refinement can degrade performance, as models often struggle to assess the correctness of their own reasoning.
Recently, inspired by the Society of Mind philosophy, Multi-Agent Debate (MAD; Chan et al. 2023; Du et al. 2023; Liang et al. 2023) has emerged as a promising alternative. However, its success has been limited to static problem-solving and lacks assessments for adversarial strategic reasoning (Cobbe et al., 2021; Edwards, 1994; He et al., 2020). Additionally, current evaluations do not account for dynamic decision-making in interactive environments, where an agent’s choices influence and adapt to an opponent’s actions. This limitation hinders LLMs from retrieving and applying strategic knowledge beyond the given context.
Table 1: Feature comparison between self-correction methods. In contrast to Multi-Agent Debate, DReaMAD encourages diverse viewpoints by varying prompts and enhancing debate robustness through automated knowledge structuring.
Methods Single Model Usage Multiple Instances Rethinking Process Self-Feedback Diversity in reasoning Perspective Shifts
Self-Consistency
Self-Refinement
Multi-Agent Debate
DReaMAD (Ours)
: means diversity from LLM’s randomness, controlled by temperature hyperparameter.
To overcome the above limitation, we introduce MetaNIM Arena, a framework for evaluating LLMs in adversarial strategic decision-making. It allows us to assess their ability to adapt dynamically and ensures robust reasoning under mathematically rigorous conditions. Through MetaNIM Arena, we systematically analyze two fundamental limitations of MAD:
@@
    Bias Reinforcement: In strategic reasoning tasks, LLMs tend to rely on immediate context rather than retrieving broader strategic knowledge, leading to distorted reasoning instead of correct inference. Debate-based frameworks further amplify this issue by reinforcing the model’s inherent biases rather than mitigating them (§4.1-4.2).
@@
    Lack of Perspective Diversity: Although MAD uses a debate structure, it relies on multiple instances of the same model. This limits the diversity of perspectives introduced in the reasoning process, reducing its ability to challenge inherent biases (§4.3).
@@
Rooted in the Learning from Multiple Approaches framework (Council et al., 2005; Cleaves, 2008), research shows that engaging with multiple problem-solving representations enhances comprehension and mitigates biases. Building on this insight, we propose DReaMAD (Diverse Reasoning via Multi-Agent Debate with Refined Prompt). DReaMAD addresses the limitations of MAD by (1) refining LLMs’ domain-specific knowledge to guide more accurate strategic reasoning, and (2) systematically modifying prompts to foster diverse perspectives. A detailed comparison with existing self-correction methods is presented in Table 1. Our key contributions are as follows:
  • We introduce MetaNIM Arena, a benchmark designed to evaluate LLMs in adversarial strategic decision-making, where mathematical rigor enables precise assessment of reasoning quality and strategic adaptability.
  • We identify the bias reinforcement in Multi-Agent Debate, showing that MAD strengthens both correct and incorrect reasoning rather than inherently improving it.
@@
  • We propose DReaMAD, a novel framework that refines strategic prior knowledge and enhances reasoning diversity through structured self-prompt refinement and perspective diversification, achieving a +12.0% accuracy gain over standard prompting on MetaNIM Arena dataset and a +20.8% higher win rate than MAD in the simulator.
2 Preliminary
2.1 Bias in LLMs
Large Language Models (LLMs) can exhibit biases that lead to unfair or skewed outcomes, arising from training data, model architectures, learning objectives, or deployment conditions (Guo et al., 2024). Such biases manifest both intrinsically, for instance in word embeddings (Bolukbasi et al., 2016), and extrinsically, reflecting real-world disparities (Goldfarb-Tarrant et al., 2021). Moreover, biases can emerge dynamically during interactive reasoning, where current reinforcement mechanisms—like self-consistency and self-refinement—often fail to mitigate them (Huang et al., 2024; Shin et al., 2024). Indeed, recent research shows that iterative interactions can reinforce existing biases instead of diversifying reasoning (Ganguli et al., 2023).
2.2 Prompt Engineering and Self-Correction in LLMs
Prompt engineering shapes model outputs without retraining, potentially improving generalization and reducing bias (Brown et al., 2020; Reynolds & McDonell, 2021; Zhao et al., 2024; Schulhoff et al., 2024a; Shin et al., 2024). However, fully eliminating biases in complex reasoning remains challenging (Jiang et al., 2022; Lu et al., 2022).
Meanwhile, self-correction mechanisms in LLMs refine responses without external supervision (Ganguli et al., 2023; Liu et al., 2024; Kamoi et al., 2024). Self-consistency, for instance, ensembles multiple outputs but converges on frequent rather than correct answers (Wang et al., 2022; Chen et al., 2023c), and self-refinement can reinforce rather than fix biases (Wan et al., 2023; Shinn et al., 2023; Madaan et al., 2024; Huang et al., 2024). Feedback-loop methods such as STaR (Zelikman et al., 2022), Reflexion (Shinn et al., 2023), and SCoRe (Kumar et al., 2024) also struggle to reliably correct biases or foster diverse reasoning (Guo et al., 2024).
Figure 1: An example demonstrating how the debate process converges to a biased outcome. We observed that bias reinforcement occurs in the first debate. Blue text indicates the correct reasoning and orange text indicates the strong consistent (biased) reasoning. The second debate is omitted, as its procedure replicates the first and third; all debates use GPT-4o-mini as the debating agent.
2.3 Multi-Agent Debate in LLMs
Multi-Agent Debate (MAD) enables LLM agents to critique each other, enhancing reasoning on complex tasks (Liang et al., 2023; Du et al., 2023). ChatEval (Chan et al., 2023), a multi-agent evaluation system, simulates human judgment to assess model output quality. Optimizations include task-specific strategies for improving debate effectiveness (Smit et al., 2024) and ACC-Debate, an actor-critic framework that trains models to specialize in debates, achieving benchmark gains (Estornell et al., 2024). While these enhancements improve performance, studies reveal a key limitation: static evaluations focus on assessing predefined problems, whereas real-world decision-making often involves dynamic, interactive environments where biases can evolve. Understanding how biases shift in these settings is crucial for developing robust strategies that extend beyond conventional static benchmarks.
@@ 3 MetaNIM Arena
3.1 Overview
We introduce MetaNIM Arena, illustrated in Figure 7. It features six impartial combinatorial games, meaning both players share identical moves at each state, all information is fully observable, and each game terminates in a finite number of moves. By merging combinatorial game theory with adversarial play, MetaNIM Arena serves as a benchmark for debate-based strategic reasoning in LLMs.
Dataset and Simulator. Key game situations are systematically collected into a dataset, each accompanied by an optimal action, enabling LLMs to be tested for decision-making accuracy. These results appear in Table 3, with details in Appendix A.3. Separately, MetaNIM Arena can function as a simulator: the model encounters an adaptive opponent, so each trajectory depends on both the agent’s and the opponent’s actions. Here, binary win/loss outcomes allows evaluation by win rate. We demonstrate this approach in Table 4, with further explanation in Appendix A.4. We refer the Appendix B for details of opponent modeling.
@@ 3.1 Overview
Table 2: Bias reinforcement across models: showing that even after the debate concludes, strongly consistent actions continue to exhibit strong consistency, reinforcing biased action distributions in the Fibonacci game. A wrong bias occurs when the model’s biased response deviates from the optimal action, while a good bias refers to cases where the biased response aligns with the optimal action. (a, b) indicate the state where the remaining items are a, and player can take the items maximum to b.
GPT-4o Wrong Bias Good Bias
(20, 19) (12, 4) (7, 4) (15, 10) (16, 8) (7, 7)
Standard 0.700 0.675 0.600 0.525 0.725 0.850
+ After MAD 0.900 0.750 0.700 0.750 0.750 0.900
GPT-4o-mini Wrong Bias Good Bias
(18, 4) (12, 6) (10, 4) (15, 10) (7, 2) (15, 2)
Standard 0.700 0.875 0.600 0.975 0.950 0.975
+ After MAD 0.850 0.950 0.750 1.000 0.950 1.000
GEMINI-1.5-pro Wrong Bias Good Bias
(20, 19) (12, 4) (4, 4) (15, 10) (10, 4) (16, 8)
Standard 0.650 0.600 0.600 0.800 0.625 0.675
+ After MAD 0.700 0.650 0.800 0.800 0.800 0.700
GEMINI-1.5-flash Wrong Bias Good Bias
(12, 6) (12, 4) (7, 4) (15, 4) (20, 19) (7, 7)
Standard 0.800 0.700 0.525 0.750 0.500 0.750
+ After MAD 0.850 1.000 0.750 0.750 0.650 1.000
3.2 Why MetaNIM Arena?
MetaNIM Arena provides a rigorous environment for adversarial strategic reasoning in LLMs. Rather than isolated problem-solving or factual recall, our framework uses impartial combinatorial games, where each position’s Grundy number defines the provably correct move. We will discuss further theoretical details in the following Section 3.4. This design offers:
1. Adversarial Strategic Reasoning: Each scenario includes an opponent whose actions shape outcomes. Models must anticipate adversarial moves across multiple turns, going beyond static QA or single-step predictions. This approach tests latent strategic knowledge in an interactive, step-by-step context.
2. Clear Optimality Criterion: By the Sprague-Grundy Theorem, these games admit an optimal strategy. MetaNIM Arena thus measures how closely a model’s reasoning aligns with that strategy, instead of relying on approximate metrics like BLEU or perplexity. We also assert that MetaNIM Arena naturally supports reinforcement learning framework. By providing a binary win/loss signal and structuring gameplay as a Markov Decision Process (MDP), it enables iterative strategy refinement—an advantage often absent in static benchmarks.
3.3 Game Variants
Here, we introduce four settings of MetaNIM Arena. Detailed explanation and theoretically determined winning strategies for these games are provided in Appendix A.
  • NIM: Agents take turns removing 1 to $N$ (typically 3) objects from a set of heaps where the player who takes the last object wins. Success requires maintaining specific heap configurations to control the game’s outcome.
  • Fibonacci: A variation of NIM where an agent’s maximum removal is constrained by the opponent’s previous move. Each turn, an agent may remove between 1 and $2\times$ the opponent’s last action. This rule introduces dynamic strategy adjustments, balancing immediate gains with long-term positioning.
  • Kayles: Played with a or two row(s) of pins, where agents take turns knocking down one or two adjacent pins. The player unable to make a move loses. The challenge lies in evaluating pin configurations and predicting the opponent’s responses to optimize each turn.
  • Chomp: Played on a quadrangle grid, agents take turns consuming a “block” of chocolate along with all blocks below and to the right. The player forced to eat the top-right (or top-left) “poisoned” block loses.
  • Corner Queen: On a rectangular board, two players take turns moving a queen left, down, or diagonally down-left. The first to reach the bottom-left corner wins. Strategy lies in limiting the opponent’s options while advancing.
3.4 Combinatorial Games: Theory and Strategy
All MetaNIM Arena games are impartial combinatorial games, forming a Directed Acyclic Graph (DAG) where vertices represent game states and edges denote valid moves. The Grundy Number framework, along with the Sprague-Grundy Theorem, guarantees the existence of a winning strategy and provides a concrete method to determine it. In the MetaNIM Arena dataset, each state has a mathematically provable optimal move, allowing an LLM’s decisions to be evaluated against the theoretical optimal strategy—a key advantage for unbiased assessment.
Definition 3.1 (Grundy Number).
For a finite impartial combinatorial game under normal play (where the last player to make a valid move wins), the Grundy number (or Nimber) $G(S)$ is recursively defined as follows. If $S$ is a terminal state with no valid moves, set $G(S)=0$. Otherwise,
$G(S)\;=\;\mathrm{mex}\{\,G(S^{\prime})\;|\;S^{\prime}\!\text{ is reachable from }S\,\}\,.$
Here, $\mathrm{mex}(X)$ is the smallest nonnegative integer not in $X$.
Note that the Grundy number is well-defined for every impartial combinatorial game, since the game’s state space forms a DAG. In many combinatorial games, direct enumeration of all possible move sequences is computationally infeasible. However, with Sprague-Grundy Theorem, we can easily calculate Grundy Numbers on complex games. See Appendix A.1.1 for further discussions.
Figure 2: Bias reinforcement in NIM game by MAD. We compared initial action distribution and action distribution after 3 rounds of debates. MAD amplifies a model’s biases, making debates favor consistent but potentially incorrect responses.
4 Understanding Bias Reinforcement in Debate Process
To quantitatively analyze bias in the debate process, we define strong consistency and bias reinforcement as below.
Definition 4.1 (Strong Consistency).
Strong consistency is a model’s tendency to consistently produce the same output with high probability when given identical inputs. If a response’s probability exceeds a threshold (set at 0.5) across multiple trials, we label the behavior as strongly consistent.
This phenomenon naturally emerges in strategic decision-making contexts. As shown in Figure 1, when presented with a specific game state, the model repeatedly generates the same reasoning pattern, highlighted in orange. Rather than effectively utilizing strategic prior knowledge, the model fixates on a single line of reasoning, limiting adaptability and decision quality.
Definition 4.2 (Bias Reinforcement).
Bias reinforcement in the context of large language models refers to the phenomenon where iterative reasoning processes—such as multi-agent debates—amplify pre-existing model biases instead of mitigating them. Rather than converging toward a more accurate or optimal reasoning outcome, the debate process reinforces strongly consistent, yet potentially suboptimal or distorted, reasoning patterns.
In the rest of this section, we analyze bias reinforcement and lack of diversity in the debate process, using GPT-4o-mini and GEMINI-1.5-pro for the NIM game, and then extend the evaluation to Fibonacci with those two plus GPT-4o and GEMINI-1.5-flash.
4.1 Bias Reinforcement in MAD
In NIM: (Figure 2) We find that MAD amplifies models’ pre-existing biases rather than refining their reasoning. To investigate this, we first identify game states where each model exhibits strong consistency, i.e., consistently selecting the same action across multiple trials. For each such state, we conduct multi-agent debates using two identical agents instantiated from the same model.
Each agent generates 20 responses (40 per state) at a fixed temperature of 0.7, maintained throughout the debate. Initial action distributions (light red) are compared against post-debate distributions after three rounds (blue). If MAD functioned as a self-correction mechanism, we would expect the distribution to shift toward the optimal action.
However, we found the opposite: regardless of whether the initial reasoning was correct, the debate process consistently amplifies pre-existing biases rather than mitigating them. For instance, in Figure 2 (top-left), GPT-4o-mini initially selects a suboptimal action (Action 3) 82.5% of the time. After the debate, this frequency increases to 90.0%, while the proportion of optimal responses drops further. Rather than correcting errors, the debate reinforces strongly consistent—but incorrect—responses. This pattern persists across model families. The GEMINI models (bottom row of Figure 2) exhibit similar behavior, regardless of whether the initial bias aligns with optimal play (“good bias”) or not (“wrong bias”). In both cases, MAD strengthens the dominant trajectory without introducing new strategic insight.
Figure 1 provides a detailed view: two LLM agents receive the same input and begin debating. Despite initial divergence, they converge quickly—after the first round—on a shared line of reasoning. Crucially, this convergence occurs even when the initial consensus is incorrect, illustrating that MAD often serves as an amplifier of bias rather than a correction mechanism.
In Fibonacci: (Table 2) Extending our NIM analysis, we evaluate MAD’s effect in the Fibonacci game, a more complex setting with move constraints and dynamic interactions. As in NIM, We identify states exhibiting strong consistency and categorize them into two groups: those where consistent responses align with the optimal strategy, and those where they do not. We then apply MAD to examine how response distributions evolve post-debate.
Consistent with the NIM results, MAD reinforces the model’s initial biases in Fibonacci as well. As shown in Table 2, the frequency of initially consistent actions increases after debate, regardless of whether those actions are optimal. On average, reinforcement rises by 9.17% in GPT models and 12.29% in GEMINI models, indicating a systematic amplification of dominant reasoning patterns across architectures.
Figure 3: Decline in optimal actions over debate rounds, demonstrating the convergence toward consistently biased reasoning.
4.2 Optimal Inputs Do Not Reduce Bias in MAD
Figure 4: DReaMAD framework. DReaMAD improves LLM reasoning by combining Strategic Prior Knowledge Elicitation and Perspective Diversification. In the first stage, the model reinterprets the problem and formulates high-level strategies to reduce bias. In the second stage, multiple agents adopt distinct viewpoints, engage in structured debate, and refine their conclusions to enhance decision-making.
To further assess the bias reinforcement in MAD, we conduct an experiment on the NIM game (5 items remaining) using GPT-4o-mini, illustrated in Figure 3. We prepare two sets of 20 responses: one exhibiting strong consistency toward a biased action (Action 2), and another consisting of optimal responses, where 80% of actions correspond to the game-theoretic best move (Action 1). We then introduced these responses into a multi-agent debate setting using GPT-4o-mini, pairing each strongly consistent response with an optimal response, and observed how the model’s reasoning evolves over multiple rounds of debate (denoted the game situation in detail in Appendix  E).
Surprisingly, the debate fails to leverage the high-quality input. Initially, the curated dataset contained 80% optimal responses, yet after a single round of debate, the model predominantly aligned with the biased responses. As debate rounds proceed, this effect intensifies: the influence of the optimal input diminishes, while the initially frequent—but suboptimal—choice becomes dominant. These results demonstrate that MAD is not only ineffective at correcting bias, but can systematically aligns with its pre-existing consistency patterns, reinforcing suboptimal but frequent choices.
In contrast, applying DReaMAD with the curated optimal responses preserves correct reasoning and mitigating bias reinforcement. This underscores the need for mechanisms that introduce diverse perspectives beyond internal debate, which we further discuss in §5.
4.3 Lack of Reasoning Diversity in MAD
While MAD is designed to refine reasoning through agent interaction, its effectiveness is fundamentally constrained by a lack of genuine diversity (§2.3). To address this limitation, Chen et al. (2023b) propose a multi-model debate framework that combines outputs from different models to encourage diverse reasoning. In contrast, our approach focuses on single-model self-correction.
Interestingly, we observe that even within a single model, small variations in prompts can induce markedly different reasoning paths (Appendix D). For example, including or omitting the word Fibonacci leads to distinct strategies in the same task. This suggests that diversity in reasoning can be enhanced by strategically modifying prompts within the same model, providing a practical alternative to multi-model debate frameworks.
5 DReaMAD: Diverse Reasoning via MAD
Figure 5: In this example, we illustrate how the debate process converges to an optimal outcome using DReaMAD. We begin with the same current state shown in Figure 1, employing self-generated prompts for each LLM agent.
To address the limitations of Multi-Agent Debate (MAD) and improve strategic decision-making in Large Language Models (LLMs), we introduce DReaMAD (Diverse Reasoning via Multi-Agent Debate with Refined Prompt).
Our framework refines prior knowledge and ensures diverse perspectives by extending MAD in two main stages:
@@ 5 DReaMAD: Diverse Reasoning via MAD
    Strategic Prior Knowledge Elicitation: The model redefines the problem, extracts key strategic insights, and formulates a high-level strategy before reasoning.
@@
    Perspective Diversification: Multiple agents are instantiated with self-generated distinct viewpoints to engage in dialectical reasoning.
After these stages, the agents conduct a structured multi-agent debate. A final post-debate refinement step then revisits their conclusions to improve reasoning quality. The complete workflow is illustrated in Figure 4 and the detailed prompt formulation used in these two modules is well documented in the Appendix C.
[…]

### v5

# From Belief Entrenchment to Robust Reasoning in LLM Agents

@@ Abstract
Multi-Agent Debate (MAD) has emerged as a promising inference scaling method for Large Language Model (LLM) reasoning. However, it frequently suffers from belief entrenchment, where agents reinforce shared errors rather than correcting them. Going beyond merely identifying this failure, we decompose it into two distinct root causes: (1) the model’s biased static initial belief and (2) homogenized debate dynamics that amplify the majority view regardless of correctness. To address these sequentially, we propose DReaMAD (Diverse Reasoning via Multi-Agent Debate). Our framework first rectifies the static belief via strategic prior knowledge elicitation, then reshapes the debate dynamics by enforcing perspective diversity. Validated on our new MetaNIM Arena benchmark, DReaMAD significantly mitigates entrenchment, achieving a +9.5% accuracy gain over ReAct prompting and a +19.0% higher win rate than standard MAD.
[…]
@@ 1 Introduction
While Large Language Models (LLMs) have shown impressive problem-solving capabilities (Achiam et al., 2023; Dubey et al., 2024), their ability to self-correct remains inconsistent. Single-agent mechanisms like Self-Refinement (Madaan et al., 2024) or Self-Consistency (Wang et al., 2022) often degrade performance when models cannot accurately assess their own reasoning (Huang et al., 2024). To address this, Multi-Agent Debate (MAD; Chan et al. 2023; Du et al. 2023; Liang et al. 2023) has emerged as a promising paradigm for inference-time scaling, leveraging inter-agent critique to refine outputs.
Figure 1: DReaMAD framework. DReaMAD improves LLM reasoning by combining Strategic Prior Knowledge Elicitation and Perspective Diversification. First, the model reinterprets the problem and formulates high-level strategies to reduce belief entrenchment. In the second stage, multiple agents adopt distinct viewpoints, engage in structured debate, and refine their conclusions to enhance decision-making.
However, we observe that standard MAD is brittle in sequential, strategic environments, where a single suboptimal decision triggers consecutive failures that the consensus process fails to correct. Rather than challenging errors, homogeneous agents frequently converge into an “echo chamber,” amplifying inherent cognitive biases. As we demonstrate in Section 3, this failure stems from a debate dynamic that prioritizes initial response frequency over logical validity, effectively rewarding popularity rather than correctness. Unlike static tasks (Cobbe et al., 2021; Edwards, 1994; He et al., 2020) where errors can be masked, strategic games enforce a rigorous penalty for every suboptimal reasoning step, making them ideal for dissecting cascading failures.
To address these limitations, we introduce MetaNIM Arena, a rigorous benchmark for adversarial strategic decision-making. Using this environment, we identify a critical failure mode termed belief entrenchment, where interaction actually hinders error correction. We decompose this phenomenon into two root causes:
@@
    Static Initial Belief: The model’s inherent bias prior to interaction, where it relies on immediate context rather than strategic foresight.
@@
    Homogenized Debate Dynamics: A self-reinforcing mechanism where the persuasiveness of an argument is dictated by its initial popularity among homogeneous agents, suppressing valid but less probable reasoning paths.
@@
Building on the Learning from Multiple Approaches framework (Council et al., 2005), which suggests that engaging with multiple problem-solving representations mitigates bias, we propose DReaMAD (Diverse Reasoning via Multi-Agent Debate with Refined Prompt). As shown in Figure 1, DReaMAD addresses the identified root causes by (1) eliciting strategic prior knowledge to correct static belief and (2) systematically modifying prompts to foster diverse perspectives, thereby reshaping debate dynamics. Our key contributions are as follows:
  • We identify and decompose belief entrenchment in Multi-Agent Debate, showing that standard MAD acts as a dynamic amplifier of static cognitive biases.
@@
  • We propose DReaMAD, a novel framework that sequentially targets the root causes of belief entrenchment through strategic knowledge elicitation and perspective diversification. It achieves a +9.5% accuracy gain over ReAct prompting on the MetaNIM Arena dataset and a +19.0% higher win rate than MAD in the simulator.
  • We introduce MetaNIM Arena, a benchmark designed to evaluate LLMs in adversarial strategic decision-making, enabling precise assessment of reasoning robustness and strategic adaptability.
@@ 2 MetaNIM Arena
We introduce MetaNIM Arena, a benchmark designed to rigorously evaluate adversarial strategic reasoning in LLMs. Unlike static QA tasks, this benchmark requires agents to make sequential decisions in dynamic environments where optimal play is mathematically definable but non-trivial to execute.
2.1 Mathematical Foundation: Impartial Games
The arena consists of five impartial games: NIM, Fibonacci Nim, Kayles, Chomp, and Corner Queen. These games share key properties: (1) Full Observability: The entire game state is visible to both players; (2) Impartiality: Available moves depend solely on the state, not on which player is moving; (3) Termination: The game is guaranteed to end in a finite number of moves.
Figure 2: An example demonstrating how the debate process converges to a biased outcome. We observed that belief entrenchment occurs in the first debate. Blue text indicates the correct reasoning and orange text indicates the strong consistent (biased) reasoning. The second debate is omitted, as its procedure replicates the first and third; all debates use GPT-4o-mini as the debating agent.
Crucially, for any non-terminal state in these games, one player holds a provably winning strategy assuming optimal play. This is grounded in the Sprague-Grundy Theorem, which maps every impartial game state to a non-negative integer called a Grundy number (or nimber). A state with a Grundy number of 0 is a losing position (P-position), while any non-zero state is a winning position (N-position). This provides a clear, objective criterion for optimality: an agent reasoning correctly should always move the game to a state with a Grundy number of 0.
2.2 Game Variants and Complexity
MetaNIM Arena includes diverse game settings to test different aspects of strategic reasoning:
  • NIM: The foundational game where players remove objects from heaps. The winning strategy relies on the XOR sum of heap sizes (Nim-sum).
  • Fibonacci Nim: A dynamic variant where the maximum removal is constrained by the opponent’s previous move (up to $2\times$). This introduces complex state dependencies requiring forward lookahead.
  • Kayles: Players knock down pins, potentially splitting a row into two independent subgames. This tests the agent’s ability to decompose problems and analyze disjoint game components.
  • Chomp: A geometric game on a grid where players consume blocks. While a winning strategy is guaranteed to exist, no simple closed-form solution is known for general grid sizes, making it a test of heuristic search and pattern recognition.
  • Corner Queen: A spatial game equivalent to Wythoff’s game, where winning positions follow the Beatty sequence (related to the Golden Ratio), requiring precise numerical reasoning.
We provide a more detailed theoretical explanation for impartial games and each game variant in Appendix A.1.
2.3 Evaluation Modes
We utilize MetaNIM Arena in two complementary modes:
Interactive Simulator.
This mode evaluates an agent’s adaptability in a turn-based environment. The agent plays against a strong adaptive opponent powered by GPT-4o using the ReAct framework (Yao et al., 2023). To ensure rigorous evaluation, the agent is always initialized in a winning position (N-position). Performance is measured by win rate, reflecting the agent’s ability to maintain the winning advantage against an optimal or near-optimal adversary. We employ various configurations (e.g., different board sizes, starting items) as detailed in Appendix A.2.
Static Dataset.
To analyze reasoning quality at a granular level, we constructed a static dataset of critical game states. Each entry consists of a state description and its unique optimal move (calculated via Grundy numbers). This allows us to measure Decision Accuracy, defined as the percentage of turns where the agent selects the theoretically optimal move, without the noise of full game simulations.
3 Analyzing Belief Entrenchment in Debate
Before introducing our method, we systematically analyze why standard Multi-Agent Debate (MAD) often fails to correct errors and instead reinforces inherent belief. We utilize the MetaNIM Arena environment to quantitatively measure this phenomenon.
3.1 Experimental Setup and Definitions
To rigorously analyze the debate dynamics, we focus on the Fibonacci Nim game, where the optimal move is mathematically deterministic. This allows us to clearly distinguish between the model’s inherent preference and the ground-truth optimal action.
Estimating Argument Probability.
Departing from token-level log-probabilities used in prior work, we estimate the probability $p_{i}$ of an argument (action) $i$ via empirical generation frequency. Specifically, we sample $N=20$ responses at a fixed temperature ($T=0.7$) to determine the action distribution. Note that $p_{i}$ represents the model’s belief in a specific context; it is not static but shifts based on the prompting strategy.
Defining Belief and Consistency.
We conceptualize the model’s static belief as its initial action probability distribution, denoted as $p_{i}^{(0)}$, prior to any multi-agent interaction.
  • Optimal Action: The theoretically unique move that guarantees a winning strategy from the current state, serving as the ground truth for correctness.
  • Suboptimal Tendency: A systematic preference for a losing action driven by heuristic shortcuts.
  • Strong Consistency: A state where the model’s belief is highly concentrated, specifically when the dominant action’s probability exceeds the majority threshold ($p_{i}>0.5$). This indicates a solidified stance likely to dominate the consensus process.
3.2 Empirical Evidence and Modeling
@@
Figure 3: Belief entrenchment across models: showing that even after the debate concludes, strongly consistent actions continue to exhibit strong consistency, reinforcing biased action distributions in the Fibonacci game.
Figure 4: Belief Entrenchment in Action (NIM). Despite one agent proposing the optimal move, the debate converges to the suboptimal majority view, illustrating how valid reasoning is suppressed by the model’s static belief.
We conducted experiments using 3 rounds of standard MAD Du et al. (2023) with various agents against GPT-4o. Figure 3 illustrates the change in probability of the dominant action after the debate.
Observation: Belief Entrenchment.
As shown in Figure 3, actions that initially exhibited Strong Consistency ($p^{(0)}>0.5$) consistently increased their probability after debate ($p^{(3)}>p^{(0)}$), regardless of whether the action was optimal or suboptimal. Furthermore, all data points for Strong Consistency lie above the $y=x$ line. This empirically demonstrates belief entrenchment: the debate process acts as an amplifier for the majority opinion, rather than a correction mechanism based on logical validity, as illustrated in Figure 2. This leads agents to converge on a suboptimal consensus, as shown in Figure 4, where a correct counterargument is overridden by the reinforced bias, leading the agents to converge on a suboptimal consensus.
Figure 5: In this example, we illustrate how the debate process converges to an optimal outcome using our algorithm, DReaMAD. We begin with the same current state shown in Figure 2, employing self-generated prompts for each LLM agent.
Modeling the Dynamics.
To explain the mechanism behind this phenomenon, we propose a probabilistic interaction model based on pairwise agent debate. We posit that the persuasiveness of an argument is not determined by its logical validity, but is proportional to the model’s static initial belief, $p_{i}^{(0)}$. This implies that agents are more likely to accept arguments that align with their inherent biases.
Mathematically, the probability of action $i$ at round $t+1$, denoted as $p^{(t+1)}_{i}$, evolves as follows:
$p^{(t+1)}_{i}=\underbrace{{p^{(t)}_{i}}\cdot{p^{(t)}_{i}}}_{\mathclap{\text{Consensus}}}\mkern 2.0mu+\mkern 9.0mu\underbrace{2p^{(t)}_{i}(1-p^{(t)}_{i})\cdot p^{(0)}_{i}}_{\mathclap{\text{Belief-Driven Resolution}}}\,$, (1)
where the first term represents the probability of both agents agreeing on action $i$, and the second term models conflict resolution biased by the initial popularity $p_{i}^{(0)}$.
Our simple model generates the theoretical curve in Figure 3, which closely fits the empirical data points with an average residual of just $2.5\%$. This tight alignment validates our hypothesis: the debate dynamics act as an amplifier of static bias ($p^{(0)}$) rather than a filter for reasoning quality.
Decomposing the Failure.
Crucially, the strong fit of our model allows us to decompose belief entrenchment into two distinct root causes, motivating our proposed method:
@@
    Static Initial Belief ($p^{(0)}$): The model’s inherent bias prior to any interaction. If the initial probability of the correct action is low ($p^{(0)}_{\mathrm{correct}}<0.5$), the debate cannot recover it.
@@
    Homogenized Debate Dynamics: The functional dependency where an argument’s persuasiveness is dictated by its initial likelihood ($p^{(0)}$) rather than its logical validity. As evidenced by our model fit, this creates an echo chamber where homogeneous agents reinforce the dominant view regardless of its correctness.
4 DReaMAD: Counteracting Belief Entrenchment